QUESTION 7

My new processor is 84 percent faster than the old one. What is the speedup? Note that 100% faster is twice as fast.
Answer #1

Solution

For example

A is N% faster than B

Execution of B/Execution of A= 1+(N/100)

We can use the same formula to find the speedup

My new processor is 84 percent faster than the old one

Execution of New Processor/Execution of old processor= 1+(84/100)

=1+0.84

=1.84

Speedup=1.84

--

all the best